# CTP-SCU/2025011

### v1

@@ IV Examples in CFT2 and CFT4
We provide two consistency checks in this section. Consider a CFT4 living in the 4-dimensional conformally flat spacetime with the metric $\mathrm{d}s^{2}=\mathrm{d}\theta^{2}+\left(\mathrm{d}r^{2}+\sum_{i=1}^{2}\mathrm{d}x_{i}^{2}\right)/r^{2}$, where the vacuum expectation value of the stress tensor has been explicitly calculated in [14],
@@
where $a$ and $c$ are different coefficients of the trace anomaly [^3] In our notations, the trace anomaly is $\left\langle T_{\;\mu}^{\mu}\right\rangle=\frac{1}{(4\pi)^{2}}\left[c(C_{\mu\nu\rho\sigma}^{2}+\frac{2}{3}\nabla^{2}\mathcal{R})-a\mathcal{G}\right]$, with the Weyl tensor $C_{\mu\nu\rho\sigma}$, the Ricci scalar $\mathcal{R}$ and the topological Euler density $\mathcal{G}.$. Plugging Equation (20) and $D=4$ into Equation (12), the entanglement entropy between $A$ and $B$ is given by
$S(A:B)=\frac{3a+c}{4}\left(\frac{l_{2}}{l_{1}}-\frac{l_{1}}{l_{2}}\right)-\frac{3a+c}{2}\log\frac{l_{2}}{l_{1}}.$ (21)
Note that $A$, $B$ are two disjoint 3-balls with the same radius. In the 4-dimensional $\mathcal{N}=4$ $SU(N)$ super Yang-Mills gauge theory, the anomaly coefficients are [15, 16]
@@
Hence
$S(A:B)=\frac{N^{2}}{4}\left(\frac{l_{2}}{l_{1}}-\frac{l_{1}}{l_{2}}\right)-\frac{N^{2}}{2}\log\frac{l_{2}}{l_{1}}.$ (23)
Traditionally, that $A$ and $B$ are two adjacent and complementary regions is of our main interest, as depicted in the right panel of Figure 5. This case is conformally equivalent to two disjoint 3-balls with the same radius, due to the fact that any global conformal transformations are sphere-preserving [10], which has been analyzed in a more general way in Section III. Consequently, in Figure 5, the left panel and the right panel are conformally equivalent by setting that
$\frac{l_{2}}{l_{1}}\simeq\left(\frac{2R}{\epsilon}\right)^{2}-2+\mathcal{O}(\epsilon),$ (24)
if we take the adjacent limit $\epsilon\rightarrow 0$. Such a setup finally gives us the area law of entanglement entropy:
$S(A:A^{c})\simeq N^{2}\left[\left(\frac{R}{\epsilon}\right)^{2}-\log\frac{2R}{\epsilon}-2\right]+\mathcal{O}(\epsilon).$ (25)
This is actually entanglement entropy of a ball with the radius $R$, which precisely matches the holographic prediction given in [8].
@@
$S(A:A^{c})=\frac{c}{3}\log\frac{2R}{\epsilon}+\mathcal{O}(\epsilon).$
@@ V Shape independence and Ryu-Takayanagi formula
Even though we only study the spheric entangling surface in this paper, our approach is clearly independent of the specific shape of entangling regions $A$ and $B$. In other words, one can replace $A$ and $B$ with generally shaped regions, denoted as $\mathbb{X}$. The relevant entanglement entropy would then be
$S(A:B)$ $=$ $-4\pi\mathcal{E}_{\text{vac}}\mathrm{Vol}(\mathbb{X}).$ (28)
However, it should be noted that to apply this formula directly, $A$ and $B$ must be reflection-symmetric with respect to each other, as only under this condition the thermal partition function can be formulated.
To calculate the entanglement entropy between two similar, disjoint regions (same shape, different sizes) with arbitrary positions, one needs to map the configuration to the reflection-symmetric one with some conformal map. For any shape of the entangling region, there should exist a set of inversive products analogous to (13).
For adjacent scenarios, it is easy to understand that, by choosing some simple setup and then taking adjacent limit, the usual divergent entanglement entropy for arbitrarily shaped region can be obtained straightforwardly.
Manifestation of Ryu-Takayanagi formula
@@
Since any pair of similar, disjoint regions falls into the reflection-symmetric case, the entanglement entropy is proportional to the area $\mathrm{Vol}(\mathbb{X})$ defined within the hyperbolic metric $\mathrm{d}s^{2}=(\mathrm{d}r^{2}+\underset{i=1}{\overset{D-2}{\sum}}\mathrm{d}x_{i}^{2})/r^{2}.$ Note that the subregion $\mathbb{X}$ could always be understood as a ($D-1$)-dimensional minimal surface embedded in higher-dimensional AdSD+1 space. Therefore, the RT formula saying that the entanglement entropy of CFTD is proportional to the area of the codimension-2 minimal surface in the AdSD+1 spacetime has been manifested in current method.
VI Holographic entanglement entropy
Figure 6: At the time slice $t_{\text{E}}=0$, two subregions $A,B$ are living on the AdS boundary $z=0$, while the minimal cross-section $E_{W}$ (gray region) of the entanglement wedge (the region bounded by the blue tube) of $A$ and $B$ is in the AdS bulk.
It would be nice to explore the gravity dual of finite entanglement entropy in the context of AdS/CFT correspondence [18, 19, 20]. Recall that a CFT in $D$-dimensional Euclidean spacetime with the metric $\mathrm{d}s_{0}^{2}=\mathrm{d}t_{\text{E}}^{2}+\mathrm{d}y^{2}+\underset{i=1}{\overset{D-2}{\sum}}\mathrm{d}x_{i}^{2}$ is dual to a ($D+1$)-dimensional Euclidean AdS spacetime with the metric
$\mathrm{d}s^{2}=\frac{\ell_{\text{AdS}}^{2}}{z^{2}}\left(\mathrm{d}z^{2}+\mathrm{d}t_{\text{E}}^{2}+\mathrm{d}y^{2}+\sum_{i=1}^{D-2}\mathrm{d}x_{i}^{2}\right).$ (29)
The entanglement entropy between two disjoint $(D-1)$-balls $A$ and $B$ with the same radius is dual to a ball-shaped region at the time slice $t_{\text{E}}=0$ of the AdSD+1 bulk, i.e., the entanglement wedge cross-section,
$E_{W}:=\left\{(x,z)|\left(z-\frac{l_{2}+l_{1}}{2}\right)^{2}+\sum_{i=1}^{D-2}x_{i}^{2}<\left(\frac{l_{2}-l_{1}}{2}\right)^{2}\right\}.$ (30)
The area of the entanglement wedge cross-section is given by
$\mathsf{A}=\ell_{\text{AdS}}^{D-1}\pi^{\frac{D}{2}-1}\frac{\Gamma\left(\frac{D}{2}\right)}{\Gamma\left(D\right)}\left(\frac{l_{2}}{l_{1}}-1\right)^{D-1}\,_{2}F_{1}\left(D-1,\frac{D}{2};D;1-\frac{l_{2}}{l_{1}}\right).$ (31)
According to the RT formula, the corresponding entanglement entropy is
$S(A:B)=\frac{\mathsf{A}}{4G_{N}^{(D+1)}}=\frac{\ell_{\text{AdS}}^{D-1}}{4G_{N}^{(D+1)}}\pi^{\frac{D}{2}-1}\frac{\Gamma\left(\frac{D}{2}\right)}{\Gamma\left(D\right)}\left(\frac{l_{2}}{l_{1}}-1\right)^{D-1}\,_{2}F_{1}\left(D-1,\frac{D}{2};D;1-\frac{l_{2}}{l_{1}}\right),$ (32)
with the ($D+1$)-dimensional gravitational constant $G_{N}^{(D+1)}$. This coincides with our field-theoretic result (15). For instance, in $D=4$, one has
$S(A:B)$ $=$ $\frac{3a+c}{4}\left(\frac{l_{2}}{l_{1}}-\frac{l_{1}}{l_{2}}\right)-\frac{3a+c}{2}\log\frac{l_{2}}{l_{1}}$ (33)
$=$ $\frac{\pi\ell_{\text{AdS}}^{3}}{8G_{N}^{(5)}}\left(\frac{l_{2}}{l_{1}}-\frac{l_{1}}{l_{2}}-2\log\frac{l_{2}}{l_{1}}\right).$ (34)
with two central charges $a=c=\pi\ell_{\text{AdS}}^{3}/8G_{N}^{(5)}$ [21, 22].
It is straightforward to extend the entanglement wedge cross-section to the general case, such as the cavity configuration we have addressed in the CFT side. This is because the AdS bulk and the dual CFT always share the same conformal symmetry. The entanglement wedge cross-section in the general configuration can be conformally mapped onto the reflection-symmetric case studied above.
@@ VI Holographic entanglement entropy
VII Conclusion

### v2

@@ IV Examples in CFT2 and CFT4
We provide two checks in this section. Consider a CFT4 living in the 4-dimensional conformally flat spacetime with the metric $\mathrm{d}s^{2}=\mathrm{d}\theta^{2}+\left(\mathrm{d}r^{2}+\sum_{i=1}^{2}\mathrm{d}x_{i}^{2}\right)/r^{2}$, where the vacuum expectation value of the stress tensor has been explicitly calculated in [14],
@@
where $a$ and $c$ are different coefficients of the trace anomaly [^2] In our notations, the trace anomaly is $\left\langle T_{\;\mu}^{\mu}\right\rangle=\frac{1}{(4\pi)^{2}}\left[c(C_{\mu\nu\rho\sigma}^{2}+\frac{2}{3}\nabla^{2}\mathcal{R})-a\mathcal{G}\right]$, with the Weyl tensor $C_{\mu\nu\rho\sigma}$, the Ricci scalar $\mathcal{R}$ and the topological Euler density $\mathcal{G}.$. In the 4-dimensional $\mathcal{N}=4$ $SU(N)$ super Yang-Mills gauge theory, the anomaly coefficients are [15, 16]
@@
Hence from equation (15), for the juxtaposed configuration,
$S_{\text{disj}}(A:B)=\frac{N^{2}}{4}\left(\frac{l_{2}}{l_{1}}-\frac{l_{1}}{l_{2}}\right)-\frac{N^{2}}{2}\log\frac{l_{2}}{l_{1}}.$ (21)
For the right panel of Figure 6, from equation (18), the adjacent entanglement entropy $S_{\text{adj}}(A:A^{c})$ of a 33-ball in $\mathbb{R}^{3}$ is
$S_{\text{adj}}(A:A^{c})\simeq N^{2}\left[\left(\frac{R}{\epsilon}\right)^{2}-\log\frac{2R}{\epsilon}-2\right]+\mathcal{O}(\epsilon),$ (22)
which precisely matches the holographic prediction given in [8].
@@
$S_{\text{adj}}(A:A^{c})=\frac{c}{3}\log\frac{2R}{\epsilon}+\mathcal{O}(\epsilon).$
@@ V.1 Shape independence
Although we only study the spherical entangling surface in this paper, our approach is clearly independent of the specific shape of entangling regions $A$ and $B$ in the reflection-symmetric juxtaposed configuration. One simply replaces $A$ and $B$ with a generally shaped region, denoted as $\mathbb{X}$. The entanglement entropy would then be
$S_{\text{disj}}(A:B)$ $=$ $-8\pi\mathcal{E}_{\text{vac}}\mathrm{Vol}(\mathbb{X}),$ (25)
where $\mathrm{Vol}(\mathbb{X})$ is easily calculated with the conformal metric (3).
A particular interesting object is the divergent adjacent entanglement entropy $S_{\text{adj}}(\mathbb{Y}:\mathbb{Y}^{c})$ for an arbitrarily shaped region $\mathbb{Y}$ in $\mathbb{R}^{D}$. It can be calculated as follows:
  • We start with $S_{\text{disj}}(\mathbb{Y}_{1}:\mathbb{Y}_{2})$ in the cavity configuration. $\mathbb{Y}_{1}$ and $\mathbb{Y}_{2}$ are nested, centered at the same point, different in scale.
  • To calculate $S_{\text{disj}}(\mathbb{Y}_{1}:\mathbb{Y}_{2})$, we need to map it to the juxtaposed configuration.
  • Typically we end up with a non-reflection-symmetric juxtaposed configuration.
  • Since the thermal partition function usually cannot be formulated for generic non-reflection-symmetric juxtaposed configurations, analytic expression of $S_{\text{disj}}(\mathbb{Y}_{1}:\mathbb{Y}_{2})$ may not be available. In this case, we can compute the partition function on the non-reflection-symmetric torus numerically.
  • Finally, take the adjacent limit,
    $S_{\text{adj}}(\mathbb{Y}:\mathbb{Y}^{c})=\lim_{\mathbb{Y}_{1}\to\mathbb{Y}_{2}}S_{\text{disj}}(\mathbb{Y}_{1}:\mathbb{Y}_{2}).$
V.2 Manifestation of Ryu-Takayanagi formula
@@
Since any pair of similar, disjoint regions falls into the reflection-symmetric case, the entanglement entropy is proportional to the volume $\mathrm{Vol}(\mathbb{X})$ defined within the hyperbolic metric $\mathrm{d}s^{2}=(\mathrm{d}r^{2}+\underset{i=1}{\overset{D-2}{\sum}}\mathrm{d}x_{i}^{2})/r^{2}.$ Note that the subregion $\mathbb{X}$ could always be understood as a (D1D-1)-dimensional minimal surface embedded in AdSD+1 space, up to a conformal transformation for asymmetric configurations. Therefore, the RT formula saying that the entanglement entropy of CFTD is proportional to the area of the codimension-2 minimal surface in the AdSD+1 spacetime has been manifested in current method.
@@
It is very interesting to note that, the disjoint entanglement entropy $S_{\text{disj}}(A:B)$ is proportional to the scaled volume of the entangling region., rather than the area of the entangling surface. As taking the adjacent limit, the contribution from the entangling surface dominates and the area law emerges.
VI Conclusion
@@
In [18], some parts of finite contributions to the adjacent entanglement entropy were extracted for $D>2$ under specific limits, including: (i) entropy between the interior and exterior of a spatial domain for a massive scalar field; (ii) numerically computed entropy for an interval in CFT. Extending our framework to massive QFT is a critical research avenue. Additionally, [19] established the universal logarithmic term of adjacent entanglement entropy in even-dimensional CFT. A particularly intriguing direction is to recover the hyperbolic mapping introduced in that work from the adjacent limits of our results.